51.5g of aqueous phosphoric acid is mixed with 52.0g of aqueous sodium hydroxide in a calorimeter. The initial temperature was 22.3 Celsius (C) and the final temperature was 31.5 Celsius (C). The calorimeter constant is 31.72 J/C and the heat capacity of the final system is 4.00 J/gC. Based on this data the enthalpy of the reaction (Hrxn) would be calculated as ________ kJ(report to 3 sig figs and include appropriate sign).
Answer #1

Total mass of solution = mass of phosphoric acid + mass of sodium hydroxide = 51.5 + 52 g = 103.5 g

Change in temperature = (final temperature - initial temperature) = (31.5 - 22.3) 0 C = 9.2 0 C

Heat gained by reaction mixture (H1) = mass of mixture × heat capacity of mixture × temperature change = 103.5 g × 4.0 J/( g 0 C) × 9.2 0 C = 3808.8 J

Heat gained by calorimeter (H2) = calorimeter constant × temperature change = 31.72 J/ C × 9.2 C = 291.8 J

Reaction enthalpy = Total heat gained by system = 3808.8 + 291.8 J = 4100.6 J = 4.1 kJ

Enthalpy of reaction is 4.1 kJ